# A Structural Characterization of the Hit Image in the Motivic Steenrod Algebra

### v3

@@ 5.2. Extension of Kameko’s non-hit classes
We now show that Kameko’s non-hit monomials and our parity refinements extend to arbitrary algebraically closed fields of characteristic 0.
@@
Suppose $z_{k}\in A^{>0,*}\cdot M_{n}(K)$. Applying $\Phi_{K}$, we obtain $\Phi_{K}(z_{k})\in A^{>0,*}\cdot M_{n}(\mathbb{C})$. But by Proposition 2.3, $\Phi(z_{k})$ is not in this image. This is a contradiction. Hence $z_{k}$ is not hit in $M_{n}(K)$. ∎
The same argument applies to any linear combination of the translates $\sigma(z_{k})$.
@@
Under these identifications, the definitions of the subspaces $H_{n}$, $G_{n}$ and the sets $M_{0},M_{1}$ (hence also the subspace $V$ and the parity functional $\varepsilon$) are preserved. Moreover, the $M_{1}$–projection $\theta$ from Definition 3.6 is functorial with respect to $\Phi_{K}^{\sharp}$ (it is defined using the same direct-sum splitting and the same basis $\pi(M_{0})\cup\pi(M_{1})$), hence $\overline{\Phi}_{K}$ carries odd-parity classes to odd-parity classes.
Therefore, if $u=\sum_{\sigma\in S}\sigma(z_{k})$ has odd parity, then its image $\overline{\Phi}_{K}(\bar{u})\in QN_{n}^{d,*}(\mathbb{C})$ is an odd-parity class, which is non-zero by Theorem 3.8 (over $\mathbb{C}$). Since $\overline{\Phi}_{K}$ is an isomorphism, $\bar{u}\neq 0$ in $QN_{n}^{d,*}(K)$. Finally, using $QN_{n}^{d,*}(K)\cong QM_{n}^{d,*}(K)/\tau\,QM_{n}^{d,*}(K)$, we conclude that $u$ represents a non-zero class in $QM_{n}^{d,*}(K)$ as well. ∎
@@ 5.3. Motivic Peterson counterexamples over $K$
We now combine the arithmetic from Section 4 with the parity statement over general base fields.
@@
The inequality $\beta(d)>n$ is purely numerical and independent of the base field; it is proved in Proposition 4.2.
For the non-vanishing of $QM_{n}^{d,*}(K)$, we may argue as follows. Over $\mathbb{C}$, Theorem 1.3 shows that $z_{k}$ is non-hit and has degree $d$. Hence $QM_{n}^{d,*}(\mathbb{C})\neq 0$. The isomorphism $M_{n}(K)\cong M_{n}(\mathbb{C})$ of $A^{*,*}$–modules induces an isomorphism
$QM_{n}^{d,*}(K)\cong QM_{n}^{d,*}(\mathbb{C}),$
so $QM_{n}^{d,*}(K)\neq 0$ as well. Alternatively, one can invoke Proposition 5.2, which implies that $z_{k}$ represents a non-zero class in $QM_{n}^{d,*}(K)$. ∎
We now state and prove the analogue for the new family of degrees from Theorem 4.4.
@@
Thus these degrees $(n,d)$ form an infinite family of counterexamples to the motivic Peterson-type conjecture over $K$, distinct from the family with k=n3k=n-3.
@@
The inequality $\beta(d)>n$ follows from Theorem 4.4 and Proposition 2.1; this is purely number-theoretic and independent of $K$.
For the non-vanishing of $QM_{n}^{d,*}(K)$ and the explicit non-hit elements, we use Proposition 5.3. In the present situation, the degree $d$ is again of the form $k+2d_{1}$ with $d_{1}=(n-1)(2^{k}-1)$, so the odd-parity non-hit theorem applies. In particular, any $u_{S}$ with $|S|$ odd has odd parity in the $M_{1}$–summand and hence defines a non-zero class in $QN_{n}^{d,*}(K)$ and therefore in $QM_{n}^{d,*}(K)$.
In particular, taking $S=\{\mathrm{id}\}$ shows that $z_{k}$ itself is non-hit and yields $QM_{n}^{d,*}(K)\neq 0$. This completes the proof. ∎
Combining Theorems 5.4 and 5.5 proves Theorem 1.7 from the introduction.

### v4

@@ 4.2. Extension of Kameko’s non-hit classes
We now show that Kameko’s non-hit monomials and our local parity criterion extend to arbitrary algebraically closed fields of characteristic 0.
@@
Suppose $z_{k}\in A^{>0,*}\cdot M_{n}(K)$. Applying $\Phi_{K}$, we obtain $\Phi_{K}(z_{k})\in A^{>0,*}\cdot M_{n}(\mathbb{C})$. But by Proposition 2.4, $\Phi_{K}(z_{k})$ is not in this image. This is a contradiction. Hence $z_{k}$ is not hit in $M_{n}(K)$. ∎
The same argument applies to odd-parity linear combinations of the translates $\sigma(z_{k})$.
@@
Under these identifications, the subspace $G_{n}$, the sets $M_{0},M_{1}$, the local top-layer summand $V$, the projection $\vartheta$, and the parity functional $\varepsilon$ are preserved, because they are defined by the same monomial formulas and the same weight conditions. Hence odd-parity classes over $K$ correspond exactly to odd-parity classes over $\mathbb{C}$.
Therefore, if $u=\sum_{\sigma\in S}\sigma(z_{k})$ has odd parity, then its image in $QN_{n}^{d,*}(\mathbb{C})$ is non-zero by Theorem 3.7. Since $\overline{\Phi}_{K}$ is an isomorphism, the class of $u$ is already non-zero in $QN_{n}^{d,*}(K)$. Finally, using $QN_{n}^{d,*}(K)\cong QM_{n}^{d,*}(K)/\tau\,QM_{n}^{d,*}(K)$, we conclude that $u$ represents a non-zero class in $QM_{n}^{d,*}(K)$ as well. ∎
@@ 4.3. Motivic Peterson counterexamples over $K$
We now combine Kameko’s arithmetic input in the case k=n3k=n-3 and a direct binary calculation in the case k=n4k=n-4 with the local parity statement over general base fields.
@@
Assertion (1) is Kameko’s Proposition 1.5 [3], which states that if k=n3k=n-3 and $\alpha(n-2)\geq 3$, then $\beta(d)>n$.
For the non-vanishing of $QM_{n}^{d,*}(K)$, over $\mathbb{C}$ the monomial $z_{k}$ is non-hit by Theorem 1.3, hence $QM_{n}^{d,*}(\mathbb{C})\neq 0$. Since $M_{n}(K)\cong M_{n}(\mathbb{C})$ as $A^{*,*}$–modules, it follows that $QM_{n}^{d,*}(K)\neq 0$. Equivalently, one may invoke Proposition 4.2. ∎
@@
Thus these degrees $(n,d)$ form an infinite family of counterexamples to the motivic Peterson-type conjecture over $K$. This family is distinct from the case k=n3k=n-3.
@@
We first prove $\beta(d)>n$. By Proposition 2.1, it is enough to show that $\alpha(d+n)>n$. Since k=n4k=n-4, we compute
$d+n$ $=(n-1)(2^{k+1}-2)+k+n$
$=(n-1)(2^{n-3}-2)+(n-4)+n$
$=(n-1)2^{\,n-3}-2(n-1)+2n-4$
$=(n-1)2^{\,n-3}-2.$
Now $n=2^{r}+1$, so $n-1=2^{r}$, and therefore
$d+n=2^{r}\cdot 2^{\,n-3}-2=2^{\,r+n-3}-2.$
Hence $d+n$ has binary expansion consisting of (r+n4)(r+n-4) ones followed by one zero, and thus
$\alpha(d+n)=r+n-4.$
Since $r\geq 5$, we have r+n4>nr+n-4>n, so $\alpha(d+n)>n$. Therefore $\beta(d)>n$ by Proposition 2.1.
For the non-vanishing and the explicit classes, note that
$d=k+2d_{1},\qquad d_{1}=(n-1)(2^{k}-1),$
so Theorem 3.7 applies in this degree. Hence any odd-parity sum
$u_{S}=\sum_{\sigma\in S}\sigma(z_{k})$
with $|S|$ odd represents a non-zero class in $QN_{n}^{d,*}(K)$ and therefore in $QM_{n}^{d,*}(K)$ by Proposition 4.3. Taking $S=\{\mathrm{id}\}$ gives, in particular, that $z_{k}$ itself is non-hit, so $QM_{n}^{d,*}(K)\neq 0$. ∎
Combining Theorems 4.4 and 4.5 proves Theorem 1.6 from the introduction.